### v1

# Interleaving Natural Language Prompting with Code Editing for Solving Programming Tasks with Generative AI Models

Abstract.
Nowadays, computing students often rely on both natural-language prompting and manual code editing to solve programming tasks. Yet we still lack a clear understanding of how these two modes are combined in practice, and how their usage varies with task complexity and student ability. In this paper, we investigate this through a large-scale study in an introductory programming course, collecting interactions from students during a three-day laboratory activity. Our analysis shows that students primarily use prompting to generate initial solutions, and then often enter short edit–run loops to refine their code following a failed execution. We find that manual editing becomes more frequent as task complexity increases, but most edits remain concise, with many affecting a single line of code. Higher-performing students tend to succeed using prompting alone, while lower-performing students rely more on edits. Student reflections confirm that prompting is helpful for structuring solutions, editing is effective for making targeted corrections, while both are useful for learning. These findings highlight the role of manual editing as a deliberate last-mile repair strategy, complementing prompting in AI-assisted programming workflows.
1. Introduction
Imagine a student typing the following prompt to an LLM: “Write a function that returns the position of the last occurrence of a positive number in a list”. Syntactically correct Python code is generated almost instantly. The student scans the code, manually editing the function name to match the test suite. Upon execution, several test cases fail due to a mishandled boundary condition. Instead of re-prompting the LLM, the student manually edits a logical operator. They test the function again, this time successfully.
The type of interaction just described is becoming increasingly common in programming classrooms (Vadaparty et al., 2024; Kazemitabaar et al., 2023). For many students, writing a program no longer starts with a blank editor, but with code produced by a natural language prompt. A well-crafted prompt can often produce code that serves as a good starting point, overcoming the complexities of syntax and cryptic compiler feedback that have traditionally been barriers for novices (Becker et al., 2019; Edwards et al., 2020; Denny et al., 2011). However, even for simple programming tasks, students rarely craft prompts that are successful on the very first attempt (Padurean et al., 2025). This raises the question of what students do next, and why. Do they revise their prompt, or dive into the code and make manual changes? This decision point is a critical, yet understudied, aspect of modern programming education: the role of manual code editing in an AI-assisted workflow.
Exploring this choice is important because of the concern from some educators that students may over-rely on LLMs, avoiding mastery of programming fundamentals (Prather et al., 2024b; Cambaz and Zhang, 2024). If problems appear solvable through prompting alone, students may fail to appreciate the precision and structure involved in writing and debugging code. This concern aligns with arguments highlighting limitations of natural language as a programming interface. Yellin emphasizes that natural language prompts are often insufficiently precise for specifying complex software behaviors (Yellin, 2023). Empirical evidence supports these limitations: in an upper-division software engineering course, Shah et al. found that students frequently needed to manually verify and debug code generated by GitHub Copilot (Shah et al., 2025).
In this work, we explore how students integrate these two emerging competencies – natural language prompting and manual code editing – when solving programming problems. We build on the idea of Prompt Problems, visual computational tasks that students solve via natural language prompting (Denny et al., 2024b). We use a custom platform that presents students with visually specified computational tasks, which they can iteratively solve via a combination of prompt-based code generation and direct code editing. Figure 1 presents a genuine student conversation within the platform and its corresponding state diagram. The student begins a new conversation (Start/Reset) by describing the task in natural language (Msg), then runs the LLM-generated code (RunLLM), which fails. They switch to manual editing (Edit) and run the edited code (RunEdit), but a different issue comes up. To tackle it, they attach the buggy code to a new prompt (MsgAtt), receive revised code from the LLM, and run it again (RunLLM). When that also fails, they return to manual editing. The student then iterates through three cycles of editing and running code (loop between Edit and RunEdit; only the final code is shown in the figure), and ultimately reaches a successful solution (Succ).
We collected such student interactions through the platform, and separately, asked students to reflect on their strategies after completing the task, describing what influenced their choice of when to prompt or edit and how they navigated the task. Using both these behavioral traces and student reflections, our analysis is guided by the following research questions (RQs):
RQ1: How do students interleave natural-language prompting and manual code editing while solving programming tasks?
RQ2: How does task complexity and students’ programming competence influence their code-editing behavior?
RQ3: How do students perceive the usefulness of natural-language prompting versus manual code editing when solving such tasks?
2. Related work
Generative AI Tools in Computing Education. Generative AI is reshaping computing education (Denny et al., 2024a; Prather et al., 2023). Beyond code explanation generation (MacNeil et al., 2023; Leinonen et al., 2023), buggy program repair (Zhang et al., 2024), personalized feedback, and clearer error messages (Wang et al., 2024; Phung et al., 2023), these models increasingly help with generation of educational content, including quizzes, programming tasks, and debugging exercises (Pădurean et al., 2024; Doughty et al., 2024; Ma et al., 2024; Pădurean et al., 2025; Nguyen et al., 2025). Furthermore, conversational agents based on LLMs enable interactive dialogues, simulating classroom scenarios and facilitating personalized tutoring across educational contexts (Markel et al., 2023; Nguyen et al., 2024b; Acun and Acun, 2023; Zamfirescu-Pereira et al., 2023; Schmucker et al., 2024). Despite these promising applications, research also highlights the need for careful management of AI-supported interactions to ensure that students remain engaged and can critically evaluate generated outputs, and genuinely develop computational thinking skills rather than relying passively on model-generated solutions (Kazemitabaar et al., 2024; Denny et al., 2024c; Prather et al., 2024b; Amoozadeh et al., 2024).
Prompt Problems and Prompt Programming. The concept of Prompt Problems, introduced by Denny et al. (Denny et al., 2024b), challenges students to solve computational tasks presented in a visual manner, by writing prompts for AI models. This approach has been shown to increase student engagement and lower syntactic barriers (Denny et al., 2024b). Further work built upon these tasks by supporting multilingual prompts for non-native English speakers (Prather et al., 2024a) and introducing tasks where students craft prompts to generate code equivalent to provided examples (Kerslake et al., 2024). Moreover, platforms such as Prompt Programming emerged as well, enabling iterative prompt refinements via chat-like interfaces and richer problem-solving interactions (Padurean et al., 2025). However, these studies mainly investigated scenarios where problems can be successfully solved only through natural language prompting, leaving unexplored the critical moments when students must decide between further prompting or manually editing generated code.
Natural Language Prompting vs. Code Editing. Educators raise concerns that easy access to AI-generated code can encourage passive reliance rather than active evaluation (Prather et al., 2024b; Nguyen et al., 2024a). Prior work has shown that learners may struggle to assess and refine AI-generated solutions, often accepting initial outputs without reviewing or validating them (Oli et al., 2025; Prather et al., 2024b). A similar line of research has shown that natural language can be too vague for describing exactly what a program should do as tasks get harder (Yellin, 2023; Hüttel, 2025). Studies of tools like GitHub Copilot found that students often need to manually edit code after prompting, suggesting that prompting alone usually is not enough to solve many problems (Shah et al., 2025; Vadaparty et al., 2024). However, detailed analyses investigating how students navigate prompting-to-editing transitions, and what influences their choices, remain scarce.
3. Methodology
This section describes the online web platform used for conducting our study, gives a detailed description of our classroom deployment, and offers brief descriptive statistics of the collected data.
3.1. Tool Overview
In order to conduct our study, we designed a custom online platform that promotes critical engagement with code generated by LLMs, supporting both natural-language prompting and manual code editing in a unified, interactive environment. The platform’s functionality was informed by current best practices in prompt-based programming tools (Denny et al., 2024b; Padurean et al., 2025). It is designed to foster iterative problem-solving and develop students’ critical assessment of code, debugging proficiency, and flexibility to choose between natural-language prompting and direct code editing. By allowing students to alternate between refining prompts and manually editing generated code, the platform encourages the development of programming skills beyond surface-level code generation.
The platform features a chat-based interface that supports multi-turn dialogue, allowing students to iteratively refine their prompts. For each message, we include the entire conversation history when sending it to the LLM, along with a system prompt instructing the model to exclude testing code or a main function. We employ GPT-4o mini (OpenAI, 2024) as the underlying LLM, due to its popularity and cost-effectiveness. Students can reset the conversation at any time using the “Reset” button and are required to do so upon reaching the -message limit. At any point, students can execute generated code using the “Run code” button, which evaluates their solution against a set of hidden test cases. These features reflect authentic programming workflows and encourage students to reflect on code behavior, reason through errors, and iteratively improve their solutions.
A novel key feature of the platform is the code editing functionality, as Figure 1 illustrates. For code generated in the most recent LLM response, students can select “Edit code”, which loads it into a dedicated editor, above the message input area. Here, students can review and revise chosen code manually. After editing, students can execute their modified code using the “Run code” button present in the editor, receiving feedback from the hidden test cases. Optionally, they can attach edited code to a new natural-language message and press “Send”, thus continuing the conversation with the LLM. By supporting smooth transitions between natural-language prompting and manual code editing, the platform empowers students to take greater ownership of their solutions and provides rich data for understanding how they navigate AI-assisted programming.
3.2. Study Setup
To understand how students integrate natural-language prompting and manual code editing, we deployed our platform as part of a laboratory activity in a large introductory Python programming course, taught at the University of Auckland. To help our investigation, we collected all student queries, LLM responses, code execution requests, and code edits. We logged only those edits that are executed or attached to a new prompt, ensuring that logged changes represent meaningful engagement. Students were also invited to provide open-ended reflections after finishing the programming tasks. Specifically, students were asked to comment on what they found to be the ideal combination of skills in natural-language prompting and manual code editing when solving these tasks, as well as to describe the strategies they used, including how they decided when to prompt or edit and what factors influenced their choices. As part of the course, students also completed three invigilated, computer-based code writing tests during the semester. These tests were conducted without access to an LLM and required students to write and debug code from scratch.
The laboratory activity was conducted over a three-day period towards the end of the semester. Students had access to programming problems, distributed across three difficulty categories and adapted to Python from a set of prompt problems previously used in literature for introductory programming (Padurean et al., 2025). The Basic Functions category included three single-function problems, such as counting negatives in a list, summing even numbers in a list, and finding the index of the last zero in a list. The Advanced Functions category featured more challenging single-function tasks, such as sorting a sublist within specified indices, updating a matrix by propagating s across rows and columns, and adding binary numbers represented as lists. The Classes & Functions category comprised multi-function problems requiring helper functions and the use of classes or data structures, including password validation, guiding a robot’s movement on a grid, and checking if an expression’s parentheses are balanced using a stack. The problems in the first category were required as part of the laboratory assignment, while the latter two categories were optional for students seeking additional practice.
3.3. Descriptive Statistics
A total of students engaged with the programming tasks. of them attempted the Basic Functions problems, with () successfully solving every problem. Additionally, students () attempted at least one Advanced Functions problem, and () at least one from the Classes & Functions category; students () attempted problems from all three categories. Among the students, completed all three tests throughout the semester. To analyze competence effects, we split students based on average test scores: scored and scored . The median was , which we rounded to the nearest whole number (i.e., ) to yield two groups of comparable size.
(per problem category).
(per problem category).
(per problem category).
(per problem category).
(per student cohort).
(per student cohort).
(per student cohort).
(per student cohort).
4. Results
In this section, we present results for each research question (RQ) and discuss some limitations of our work.
4.1. RQ1: Interleaving Prompts and Code Edits
Figure 2(a) presents the most frequent transitions between interaction states, aggregated across all sessions (i.e., student attempting a problem). For each session, we ordered states chronologically and counted all adjacent state pairs, then selected the top transitions by frequency to highlight the most important patterns. The figure shows that the most common interaction pattern is looping between messaging and running LLM-generated code, with a notable part transitioning directly from RunLLM to success. At the same time, there is a substantial pathway from unsuccessful execution of generated code into the Edit–RunEdit loop, indicating that many students refine solutions through manual edits and check their correctness by running them as well. Overall, the diagram suggests that both prompting and code editing are actively interleaved in students’ problem-solving workflows, though their relative use varies.
Figure 2(b) provides a turn-based flow diagram of the first nine turns per session, clarifying how students interleave prompting and editing actions over time. The diagram alternates between odd-numbered turns, which represent messaging or terminal events, and even-numbered turns, which group executions and edit actions into sets. Three insights stand out from the visualization. First, a substantial number of sessions reach a solution within the first few turns, reflecting that initial prompts can often be effective. Second, manual editing typically appears as a reaction to unsuccessful executions of LLM-generated code, as indicated by the thicker flow through the combined state {RunLLM, RunEdit}, compared to the narrower {RunEdit}-only path. This confirms students mostly introduce manual edits after testing LLM-generated code. Finally, the ratio between the RunLLM-only and RunLLM, RunEdit flows remains relatively stable across turns, indicating that students introduce edits at a wide range of points during their sessions.
4.2. RQ2: Complexity and Competence Effects
To understand how task complexity and programming ability influence students’ code-editing behavior, we analyzed editing activity by problem category and competence level. We restrict our analysis to edits that were executed as the final version before the next turn or explicitly attached to a message, aligning with the turn-based structure shown in Figure 2(b). Thus, we exclude intermediate or discarded drafts overwritten within the same turn.
Figure 3 presents four analyzed indicators in a two-row layout. The top row compares editing behavior across problem categories (Basic Functions, Advanced Functions, Classes & Functions), while the bottom row focuses on Basic Functions only, comparing three student cohorts: all students, those with an average test score (Score:High), and those with a score (Score:Low). Figures 3(a) and 3(e) show the number of problems edited per student, out of those they attempted. Figures 3(b) and 3(f) report how many edits occurred per session, normalized over all sessions in each group. Figures 3(c) and 3(g) show the number of lines changed per edit using a line-level Levenshtein distance, aggregated across all retained edits. Finally, Figures 3(d) and 3(h) present the number of additional messages sent after each edit, again measured across retained edits.
As Figures 3(a) and 3(b) show, students were more likely to edit code when working on complex problems. In contrast, editing was less frequent in Basic Functions, suggesting that prompting alone was often sufficient for simpler tasks. Because Advanced Functions and Classes & Functions were optional, we repeated the analysis on the students who attempted at least one problem in each category (visualizations for this analysis are not shown here). The trends remained consistent: editing was more common in the harder categories, with around of students editing at least one problem, compared to about in Basic Functions. Notably, the number of edited lines did not substantially change with respect to task difficulty. As shown in Figures 3(c) and 3(d), most edits involved only a single line and were followed by minimal further interaction, indicating that students used editing for small, corrective adjustments.
Figures 3(e) and 3(f) show that higher-performing students were less likely to edit code. About of students in the high-score group solved all Basic Functions problems without edits, and almost of their sessions included no edits. In comparison, the low-score group showed more frequent use of manual editing. This suggests that the high-score group was more successful using prompting alone and relied less on code editing. This is a striking result given that students in the high-score group are presumably better at code editing based on the computer-based code writing tests. One possible explanation is that stronger programmers have a better grasp of the underlying logic and structure of the problem, which helps them write clearer and more precise prompts. As a result, the LLM-generated code is more likely to be correct, reducing the need for manual corrections. As with task complexity, programming competence did not affect the size or placement of edits: Figures 3(g) and 3(h) show that most edits were brief and followed by few additional messages across cohorts.
4.3. RQ3: Student Perceptions
To understand how students perceive the roles of prompting and editing in their problem-solving process, we analyzed student reflections. Students first answered a Likert-style item asking what combination of prompting and editing they found to be most effective for these prompt-based programming tasks, choosing among five options from “only natural language prompting” to “only manual code editing”. They then responded to an open-ended question describing their own strategies of how they chose between prompting and editing, and what factors influenced their decisions.
Figure 4 summarizes the Likert-scale preferences. The most common preference was an equal mix of prompting and editing, with students perceiving the two as complementary. Open-ended reflections revealed that while prompting was valued for convenience, students also saw manual editing as essential for refining logic and fixing issues when the AI fell short. We found three key themes: prompting for speed, editing for precision, and learning with AI.
Prompting for speed. Many students emphasized the efficiency of using natural language prompts to generate code quickly, turning to editing only for small adjustments. One student explained, “If I could code it manually faster than I could prompt it I would code it manually, I use prompts only for efficiency”. Another echoed this sentiment: “I only use manual code edits when there is a simple mistake that I can fix in a matter of seconds. If fixing the mistake would take longer than 30 seconds, I write another prompt instead”. A third reflected, “I used AI prompts first to get the basic code quickly. When I needed small fixes, I edited the code myself. The AI did most of the work, and I just edited the parts that needed changes. This not only improves the speed but also ensures usability”.
Editing for precision. Several students described a hybrid workflow, where prompting served as a starting point and editing was used for fine-tuning. As one put it, “I found a mixed approach most effective: leveraging the model’s strengths for speed and foundational code generation while using manual edits to ensure accuracy, handle edge cases, and align the solution with exact requirements”. Another wrote, “I used the AI to write the general structure of the solution and then used error messages to debug and manually fix any small issues like a wrong variable name or any edge cases”. One student described: “At first I use natural language prompts and keep on editing the natural language prompts until the program executes roughly 90% of what I intended the code to be, then the rest of the 10% can be manually coded by me to be fixed”.
Learning with AI. Finally, some students reported using both modes as learning tools, choosing prompting when exploring new concepts and editing when consolidating their understanding. One wrote, “If it is a question I know how to do, then I use manual code, otherwise I use a prompt and learn from the AI generated code”. Another reflected, “I feel like I used a mix of both natural language prompts and manual code, but tried to fix the errors myself before asking for help as much as I could. This was a very cool tool to use to learn how to code as I could learn from mistakes I made, rather than just copying down the fixed code”. A third explained, “I used manual code edits because they helped me retain what I learned from my mistake. But natural language prompts are also helpful when I don’t know where to begin or can’t find out what I’ve done wrong, even after checking thoroughly. The method I used depended on how confident I felt. If I knew the issue, I’d edit the code myself”.
4.4. Limitations
We acknowledge several limitations of our study that could be tackled in future work. First, only the Basic Functions problems were mandatory, while more complex tasks were optional, introducing potential self-selection bias. Future studies could mitigate this by assigning tasks of varying difficulty to all students. Second, we did not conduct an A/B study to compare prompting-only and prompting-plus-editing conditions in a controlled way. A randomized study where editing capabilities are selectively enabled could more directly measure the benefits of code editing. Finally, our logs capture only edited code that was either executed or sent as an attachment, without information on how much time or effort students spent editing. Capturing intermediate drafts or more granular timestamps could provide deeper insights into students’ editing behavior.
5. Conclusion
Our study investigated a key question in AI-assisted programming: when do students revise their prompt, and when do they dive into the code? We found that students combined prompting and manual editing in their workflow, typically introducing edits after a generated solution failed execution. These manual edits were usually concise, serving as targeted, last-mile corrections. Edits became more common as tasks increased in complexity, though higher-performing students relied more consistently on prompting alone, reflecting their ability to craft effective prompts. Student reflections highlighted prompting for efficiency, editing for precision, and both as valuable tools for learning. These findings emphasize the need to support both skills in instruction, teaching students to adapt their approach dynamically. Developing this flexibility requires a solid understanding of the problem and the concepts involved in solving it.
We believe our work opens up exciting ideas for future research. Future work should explore how to help students craft precise prompts while building fluency in concise, high-impact edits. First, studies could introduce controlled bugs into otherwise correct completions to cultivate students’ ability to detect and correct subtle issues. Second, research might investigate how students make strategic decisions between prompting and editing by presenting tasks that require them to justify which approach would be more effective in a given context. Third, longer-term investigations could track how editing behaviors evolve in more open-ended domains such as data analysis, visualization, or machine learning, where last-mile fixes may grow into broader refactors, offering deeper insight into how students balance prompting with manual control of code.
Acknowledgements.
Funded/Cofunded by the European Union (ERC, TOPS, 101039090). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.References
- (1)
- Acun and Acun (2023) Cagla Acun and Ramazan Acun. 2023. GAI-Enhanced Assignment Framework: A Case Study on Generative AI Powered History Education. In NeurIPS’23 Workshop on Generative AI for Education.
- Amoozadeh et al. (2024) Matin Amoozadeh, Daye Nam, Daniel Prol, Ali Alfageeh, James Prather, Michael Hilton, Sruti Srinivasa Ragavan, and Amin Alipour. 2024. Student-AI Interaction: A Case Study of CS1 Students. In Proceedings of the Koli Calling International Conference on Computing Education Research (Koli Calling).
- Becker et al. (2019) Brett A. Becker, Paul Denny, Raymond Pettit, Durell Bouchard, Dennis J. Bouvier, Brian Harrington, Amir Kamil, Amey Karkare, Chris McDonald, Peter-Michael Osera, Janice L. Pearce, and James Prather. 2019. Compiler Error Messages Considered Unhelpful: The Landscape of Text-Based Programming Error Message Research. In Proceedings of the Working Group Reports of the Conference on Innovation and Technology in Computer Science Education (ITiCSE).
- Cambaz and Zhang (2024) Doga Cambaz and Xiaoling Zhang. 2024. Use of AI-driven Code Generation Models in Teaching and Learning Programming: a Systematic Literature Review. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Denny et al. (2024a) Paul Denny, Sumit Gulwani, Neil T. Heffernan, Tanja Käser, Steven Moore, Anna N. Rafferty, and Adish Singla. 2024a. Generative AI for Education (GAIED): Advances, Opportunities, and Challenges. CoRR abs/2402.01580 (2024).
- Denny et al. (2024b) Paul Denny, Juho Leinonen, James Prather, Andrew Luxton-Reilly, Thezyrie Amarouche, Brett A. Becker, and Brent N. Reeves. 2024b. Prompt Problems: A New Programming Exercise for the Generative AI Era. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Denny et al. (2011) Paul Denny, Andrew Luxton-Reilly, Ewan D. Tempero, and Jacob Hendrickx. 2011. Understanding the syntax barrier for novices. In Proceedings of the Innovation and Technology in Computer Science Education Conference (ITiCSE).
- Denny et al. (2024c) Paul Denny, Stephen MacNeil, Jaromir Savelka, Leo Porter, and Andrew Luxton-Reilly. 2024c. Desirable Characteristics for AI Teaching Assistants in Programming Education. In Proceedings Companion of the Conference on Innovation and Technology in Computer Science Education (ITiCSE).
- Doughty et al. (2024) Jacob Doughty, Zipiao Wan, Anishka Bompelli, Jubahed Qayum, Taozhi Wang, Juran Zhang, Yujia Zheng, Aidan Doyle, Pragnya Sridhar, Arav Agarwal, Christopher Bogart, Eric Keylor, Can Kültür, Jaromír Savelka, and Majd Sakr. 2024. A Comparative Study of AI-Generated (GPT-4) and Human-crafted MCQs in Programming Education. In Proceedings of the Australasian Computing Education Conference (ACE).
- Edwards et al. (2020) John Edwards, Joseph Ditton, Dragan Trninic, Hillary Swanson, Shelsey Sullivan, and Chad D. Mano. 2020. Syntax Exercises in CS1. In Proceedings of the International Computing Education Research Conference (ICER).
- Hüttel (2025) Hans Hüttel. 2025. On Program Synthesis and Large Language Models. Commun. ACM (2025).
- Kazemitabaar et al. (2023) Majeed Kazemitabaar, Xinying Hou, Austin Z. Henley, Barbara Jane Ericson, David Weintrop, and Tovi Grossman. 2023. How Novices Use LLM-based Code Generators to Solve CS1 Coding Tasks in a Self-Paced Learning Environment. In Proceedings of the Koli Calling International Conference on Computing Education Research (Koli Calling).
- Kazemitabaar et al. (2024) Majeed Kazemitabaar, Runlong Ye, Xiaoning Wang, Austin Zachary Henley, Paul Denny, Michelle Craig, and Tovi Grossman. 2024. CodeAid: Evaluating a Classroom Deployment of an LLM-based Programming Assistant that Balances Student and Educator Needs. In Proceedings of the Conference on Human Factors in Computing Systems (CHI).
- Kerslake et al. (2024) Chris Kerslake, Paul Denny, David H. Smith, James Prather, Juho Leinonen, Andrew Luxton-Reilly, and Stephen MacNeil. 2024. Integrating Natural Language Prompting Tasks in Introductory Programming Courses. In Proceedings of the Virtual Global Computing Education Conference (SIGCSE Virtual).
- Leinonen et al. (2023) Juho Leinonen, Paul Denny, Stephen MacNeil, Sami Sarsa, Seth Bernstein, Joanne Kim, Andrew Tran, and Arto Hellas. 2023. Comparing Code Explanations Created by Students and Large Language Models. In Proceedings of the Conference on Innovation and Technology in Computer Science Education (ITiCSE).
- Ma et al. (2024) Qianou Ma, Hua Shen, Kenneth Koedinger, and Sherry Tongshuang Wu. 2024. How to Teach Programming in the AI Era? Using LLMs as a Teachable Agent for Debugging. In Proceedings of the International Conference on Artificial Intelligence in Education (AIED).
- MacNeil et al. (2023) Stephen MacNeil, Andrew Tran, Arto Hellas, Joanne Kim, Sami Sarsa, Paul Denny, Seth Bernstein, and Juho Leinonen. 2023. Experiences from Using Code Explanations Generated by Large Language Models in a Web Software Development E-Book. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Markel et al. (2023) Julia M. Markel, Steven G. Opferman, James A. Landay, and Chris Piech. 2023. GPTeach: Interactive TA Training with GPT-based Students. In Proceedings of the Conference on Learning @ Scale (L@S).
- Nguyen et al. (2024b) Manh Hung Nguyen, Sebastian Tschiatschek, and Adish Singla. 2024b. Large Language Models for In-Context Student Modeling: Synthesizing Student’s Behavior in Visual Programming from One-Shot Observation. In Proceedings of the International Conference on Educational Data Mining (EDM).
- Nguyen et al. (2025) Manh Hung Nguyen, Alkis Gotovos Victor-Alexandru Pădurean, Sebastian Tschiatschek, and Adish Singla. 2025. Synthesizing High-Quality Programming Tasks with LLM-based Expert and Student Agents. CoRR abs/2504.07655 (2025).
- Nguyen et al. (2024a) Sydney Nguyen, Hannah McLean Babe, Yangtian Zi, Arjun Guha, Carolyn Jane Anderson, and Molly Q. Feldman. 2024a. How Beginning Programmers and Code LLMs (Mis)read Each Other. In Proceedings of the Conference on Human Factors in Computing Systems (CHI).
- Oli et al. (2025) Priti Oli, Rabin Banjade, Andrew M. Olney, and Vasile Rus. 2025. Can LLMs Identify Gaps and Misconceptions in Students’ Code Explanations? CoRR abs/2501.10365 (2025).
- OpenAI (2024) OpenAI. 2024. GPT-4o mini: Advancing Cost-efficient Intelligence. https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/.
- Padurean et al. (2025) Victor-Alexandru Padurean, Paul Denny, Alkis Gotovos, and Adish Singla. 2025. Prompt Programming: A Platform for Dialogue-based Computational Problem Solving with Generative AI Models. In Proceedings of the Innovation and Technology in Computer Science Education Conference (ITiCSE).
- Pădurean et al. (2025) Victor-Alexandru Pădurean, Paul Denny, and Adish Singla. 2025. BugSpotter: Automated Generation of Code Debugging Exercises. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Pădurean et al. (2024) Victor-Alexandru Pădurean, Georgios Tzannetos, and Adish Singla. 2024. Neural Task Synthesis for Visual Programming. Transactions on Machine Learning Research (TMLR) (2024).
- Phung et al. (2023) Tung Phung, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares. 2023. Generating High-Precision Feedback for Programming Syntax Errors using Large Language Models. In Proceedings of the International Conference on Educational Data Mining (EDM).
- Prather et al. (2023) James Prather, Paul Denny, Juho Leinonen, Brett A. Becker, Ibrahim Albluwi, Michelle Craig, Hieke Keuning, Natalie Kiesler, Tobias Kohn, Andrew Luxton-Reilly, Stephen MacNeil, Andrew Petersen, Raymond Pettit, Brent N. Reeves, and Jaromír Savelka. 2023. The Robots Are Here: Navigating the Generative AI Revolution in Computing Education. In Proceedings of the Working Group Reports of the Conference on Innovation and Technology in Computer Science Education (ITiCSE).
- Prather et al. (2024a) James Prather, Brent N. Reeves, Paul Denny, Juho Leinonen, Stephen MacNeil, Andrew Luxton-Reilly, João Orvalho, Amin Alipour, Ali Alfageeh, Thezyrie Amarouche, Bailey Kimmel, Jared Wright, Musa Blake, and Gweneth Barbre. 2024a. Breaking the Programming Language Barrier: Multilingual Prompting to Empower Non-Native English Learners. CoRR abs/2412.12800 (2024).
- Prather et al. (2024b) James Prather, Brent N. Reeves, Juho Leinonen, Stephen MacNeil, Arisoa S. Randrianasolo, Brett A. Becker, Bailey Kimmel, Jared Wright, and Ben Briggs. 2024b. The Widening Gap: The Benefits and Harms of Generative AI for Novice Programmers. In Proceedings of the Conference on International Computing Education Research (ICER).
- Schmucker et al. (2024) Robin Schmucker, Meng Xia, Amos Azaria, and Tom M. Mitchell. 2024. Ruffle &Riley: Insights from Designing and Evaluating a Large Language Model-Based Conversational Tutoring System. In Proceedings of the International Conference on Artificial Intelligence in Education (AIED).
- Shah et al. (2025) Anshul Shah, Anya Chernova, Elena Tomson, Leo Porter, William G. Griswold, and Adalbert Gerald Soosai Raj. 2025. Students’ Use of GitHub Copilot for Working with Large Code Bases. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Vadaparty et al. (2024) Annapurna Vadaparty, Daniel Zingaro, David H. Smith IV, Mounika Padala, Christine Alvarado, Jamie Gorson Benario, and Leo Porter. 2024. CS1-LLM: Integrating LLMs into CS1 Instruction. In Proceedings of the Innovation and Technology in Computer Science Education Conference (ITiCSE).
- Wang et al. (2024) Sierra Wang, John C. Mitchell, and Chris Piech. 2024. A Large Scale RCT on Effective Error Messages in CS1. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Yellin (2023) Daniel M. Yellin. 2023. The Premature Obituary of Programming. Commun. ACM (2023).
- Zamfirescu-Pereira et al. (2023) J.D. Zamfirescu-Pereira, Laryn Qi, Bjorn Hartmann, John DeNero, and Narges Norouzi. 2023. Conversational Programming with LLM-Powered Interactive Support in an Introductory Computer Science Course. In NeurIPS’23 Workshop on Generative AI for Education.
- Zhang et al. (2024) Jialu Zhang, José Pablo Cambronero, Sumit Gulwani, Vu Le, Ruzica Piskac, Gustavo Soares, and Gust Verbruggen. 2024. PyDex: Repairing Bugs in Introductory Python Assignments using LLMs. Proceedings of the ACM on Programming Languages (2024).