### v2

# Interleaving Natural Language Prompting with Code Editing for Solving Programming Tasks with Generative AI Models

Abstract.
Modern computing students often rely on both natural-language prompting and manual code editing to solve programming tasks. Yet we still lack a clear understanding of how these two modes are combined in practice, and how their usage varies with task complexity and student ability. In this paper, we investigate this through a large-scale study in an introductory programming course, collecting interactions from students during a three-day lab activity. Our analysis shows that students primarily use prompting to generate initial solutions, and then often enter short edit-run loops to refine their code following a failed execution. Student reflections confirm that prompting is helpful for structuring solutions, editing is effective for making targeted corrections, while both are useful for learning. We find that manual editing becomes more frequent as task complexity increases, but most edits remain concise, with many affecting a single line of code. Higher-performing students tend to succeed using prompting alone, while lower-performing students rely more on edits. These findings highlight the role of manual editing as a deliberate last-mile repair strategy, complementing prompting in AI-assisted programming workflows.
1. Introduction
Imagine a student typing the following prompt to a generative AI (GenAI) model: “Write a function returning the position of the last occurrence of a positive number in a list”. Syntactically correct Python code is generated almost instantly. The student scans the code, manually editing the function name to match the test suite. Upon execution, several test cases fail due to a mishandled boundary condition. Instead of re-prompting the GenAI, the student manually edits a logical operator. They test the function again, this time successfully.
The type of interaction just described is becoming increasingly common in programming classrooms (Vadaparty et al., 2024; Kazemitabaar et al., 2023). For many students, writing a program no longer starts with a blank editor, but with code produced by a natural language prompt. A well-crafted prompt can often produce code that serves as a good starting point, overcoming the complexities of syntax and cryptic compiler feedback that have traditionally been barriers for novices (Becker et al., 2019; Edwards et al., 2020; Denny et al., 2011). However, even for simple programming tasks, students rarely craft prompts that are successful on the very first attempt (Padurean et al., 2025). This raises the question of what students do next, and why. Do they revise their prompt, or dive into the code and make manual changes? This decision point is a critical, yet understudied, aspect of modern programming education: the role of manual code editing in an AI-assisted workflow.
Exploring this choice is important because of the concern from some educators that students may over-rely on GenAI, avoiding mastery of programming fundamentals (Prather et al., 2024b; Cambaz and Zhang, 2024). If problems appear solvable through prompting alone, students may fail to appreciate the precision and structure involved in writing and debugging code. This concern aligns with arguments highlighting limitations of natural language as a programming interface. Yellin emphasizes that natural language prompts are often insufficiently precise for specifying complex software behaviors (Yellin, 2023). Empirical evidence supports these limitations: in an upper-division software engineering course, Shah et al. found that students frequently needed to manually verify and debug code generated by GitHub Copilot (Shah et al., 2025).
In this work, we explore how students integrate these two emerging competencies – natural language prompting and manual code editing – when solving programming problems. We build on the idea of Prompt Problems, visual computational tasks that students solve via natural language prompting (Denny et al., 2024b). We use a publicly available web-based platform that presents students with visually specified computational tasks, which they can iteratively solve via a combination of prompt-based code generation and direct code editing. Figure 1 presents a genuine student conversation within the platform and its corresponding state diagram. The student begins a new conversation (Start/Reset) by describing the task in natural language (Msg), then runs the GenAI-generated code (RunAI), which fails. They switch to manual editing (Edit) and run the edited code (RunEdit), but a different issue comes up. To tackle it, they attach the buggy code to a new prompt (MsgAtt), receive revised code from the GenAI, and run it again (RunAI). When that also fails, they return to manual editing. The student then iterates through three cycles of editing and running code (loop between Edit and RunEdit; only the final code is shown in the figure), and ultimately reaches a successful solution (Succ).
We collected such student interactions through the platform, and separately, asked students to reflect on their strategies after completing the task, describing what influenced their choice of when to prompt or edit and how they navigated the task. Using both these behavioral traces and student reflections, our analysis is guided by the following research questions (RQs):
RQ1: How do students interleave natural-language prompting and manual code editing while solving programming tasks w.r.t. both behavioral patterns and perceived usefulness?
RQ2: How does task complexity influence students’ code-editing behavior?
RQ3: How does students’ programming competence influence their code-editing behavior?
2. Related Work
Generative AI Tools in Computing Education. Generative AI is reshaping computing education (Denny et al., 2024a; Prather et al., 2023). Beyond code explanation generation (MacNeil et al., 2023; Leinonen et al., 2023), buggy program repair (Zhang et al., 2024), personalized feedback, and clearer error messages (Wang et al., 2024; Phung et al., 2023), these models increasingly help with generation of educational content, including quizzes, programming tasks, and debugging exercises (Pădurean et al., 2024; Doughty et al., 2024; Ma et al., 2024; Pădurean et al., 2025; Nguyen et al., 2025). Furthermore, conversational agents based on GenAI enable interactive dialogues, simulating classroom scenarios and facilitating personalized tutoring across educational contexts (Markel et al., 2023; Nguyen et al., 2024b; Acun and Acun, 2023; Zamfirescu-Pereira et al., 2023; Schmucker et al., 2024). Despite these promising applications, research also highlights the need for careful management of AI-supported interactions to ensure that students remain engaged and can critically evaluate generated outputs, and genuinely develop computational thinking skills rather than relying passively on model-generated solutions (Kazemitabaar et al., 2024; Denny et al., 2024c; Prather et al., 2024b; Amoozadeh et al., 2024).
Prompt Problems and Prompt Programming. The concept of Prompt Problems, introduced by Denny et al. (Denny et al., 2024b), challenges students to solve computational tasks presented in a visual manner, by writing prompts for AI models. This approach has been shown to increase student engagement and lower syntactic barriers (Denny et al., 2024b). Further work built upon these tasks by supporting multilingual prompts for non-native English speakers (Prather et al., 2024a) and introducing tasks where students craft prompts to generate code equivalent to provided examples (Kerslake et al., 2024). Moreover, platforms such as Prompt Programming emerged as well, enabling iterative prompt refinements via chat-like interfaces and richer problem-solving interactions (Padurean et al., 2025). However, these studies mainly investigated scenarios where problems can be successfully solved only through natural language prompting, leaving unexplored the critical moments when students must decide between further prompting or manually editing generated code.
Natural Language Prompting vs. Code Editing. Educators raise concerns that easy access to AI-generated code can encourage passive reliance rather than active evaluation (Prather et al., 2024b; Nguyen et al., 2024a). Prior work has shown that learners may struggle to assess and refine AI-generated solutions, often accepting initial outputs without reviewing or validating them (Oli et al., 2025; Prather et al., 2024b). A similar line of research has shown that natural language can be too vague for describing exactly what a program should do as tasks get harder (Yellin, 2023; Hüttel, 2025). Studies of tools like GitHub Copilot found that students often need to manually edit code after prompting, suggesting that prompting alone usually is not enough to solve many problems (Shah et al., 2025; Vadaparty et al., 2024). However, detailed analyses investigating how students navigate prompting-to-editing transitions, and what influences their choices, remain scarce.
3. Methodology
This section describes the online web platform used for conducting our study, gives a detailed description of our classroom deployment, and offers brief descriptive statistics of the collected data.
3.1. Tool Overview
To conduct our study, we used the Prompt Programming platform (Padurean et al., 2025). The publicly available platform supports a chat-based workflow where students can solve programming problems by writing natural-language prompts, manually editing generated code, and running solutions against hidden tests. We selected our problems from the platform’s existing problem set, and contacted the platform maintainers to obtain anonymized interaction logs for our analysis. We provide further details about the platform next.
The platform features a chat-based interface that supports multi-turn dialogue, allowing students to iteratively refine their prompts. For each message, the full conversation history is provided to the GenAI along with a system instruction to exclude testing code and a main function. We employ GPT-4o mini (OpenAI, 2024) as the underlying GenAI model, due to its popularity and cost-effectiveness. Students can reset the conversation at any time using the “Reset” button and are required to do so upon reaching the -message limit. At any point, students can execute generated code using the “Run code” button, which evaluates their solution against a set of hidden test cases. These features reflect authentic programming workflows and encourage students to reflect on code behavior, reason through errors, and iteratively improve their solutions.
A key feature of the platform is the code editing functionality, as Figure 1 illustrates. For code generated in the most recent GenAI response, students can select “Edit code”, which loads it into a dedicated editor, above the message input area. Here, students can review and revise chosen code manually. After editing, students can execute their modified code using the “Run code” button present in the editor, receiving feedback from hidden test cases. Optionally, they can attach edited code to a new natural-language message and press “Send”, thus continuing the conversation with the GenAI. By supporting smooth transitions between natural-language prompting and manual code editing, the platform empowers students to take greater ownership of their solutions and provides rich data for understanding how they navigate AI-assisted programming.
3.2. Study Setup
To understand how students integrate natural-language prompting and manual code editing, we used the study platform as part of a laboratory activity in a large introductory Python programming course, taught at the University of Auckland. To help our investigation, we collected all student queries, GenAI responses, code execution requests, and code edits. We logged only those edits that are executed or attached to a new prompt, ensuring that logged changes represent meaningful engagement. Students were also invited to provide open-ended reflections after finishing all the programming tasks. Specifically, students were asked to comment on what they found to be the ideal combination of skills in natural-language prompting and manual code editing when solving these tasks, as well as to describe the strategies they used, including how they decided when to prompt or edit and what factors influenced their choices. As part of the course, students also completed three invigilated, computer-based code writing tests during the semester. These tests took place without access to a GenAI model and required students to write and debug code from scratch.
The laboratory activity was conducted over a three-day period towards the end of the semester. Students had access to programming problems, distributed across three difficulty categories and adapted to Python from a set of prompt problems previously used in literature for introductory programming (Padurean et al., 2025). The Basic Functions category included three single-function problems, referred to as P1, P2, and P3 in our later analysis: counting negatives in a list, summing even numbers in a list, and finding the index of the last zero in a list, respectively. The Advanced Functions category featured more challenging single-function tasks, such as sorting a sublist within specified indices, updating a matrix by propagating s across rows and columns, and adding binary numbers represented as lists. The Classes & Functions category comprised multi-function problems requiring helper functions and the use of classes or data structures, including password validation, guiding a robot’s movement on a grid, and checking if an expression’s parentheses are balanced using a stack. The problems in the first category (P1, P2, and P3) were required as part of the laboratory assignment, while the latter two categories were optional for students seeking additional practice.
3.3. Descriptive Statistics and Analysis
A total of students engaged with the programming tasks. of them attempted the Basic Functions problems, with () successfully solving every problem. Additionally, students () attempted at least one Advanced Functions problem, and () at least one from the Classes & Functions category; students () attempted problems from all three categories. Among the students, completed all three tests throughout the semester. To analyze competence effects, we split students based on average test scores: scored and scored . The median was , which we rounded to the nearest whole number (i.e., ) to yield two groups of comparable size.
To analyze differences between problems of different difficulties and students with different competences, we conducted pairwise Mann-Whitney U tests (Mann and Whitney, 1947) to see if the distributions are different.
4. Results
In this section, we present results for each research question (RQ) and discuss some limitations of our work.
4.1. RQ1: Interleaving Prompts and Code Edits
To gain a general understanding of how students interleave natural language prompting and manual code editing, we analyze their behavioral patterns from interaction logs (see Figure 2). We then complement this with students’ perceptions from reflection responses.
Figure 2(a) presents the most frequent transitions between interaction states, aggregated across all sessions (i.e., student attempting a problem). For each session, we ordered states chronologically and counted all adjacent state pairs, then selected the top transitions by frequency to highlight the most important patterns. The figure shows that the most common interaction pattern is looping between messaging and running GenAI-generated code, with a notable subset transitioning directly from RunAI to success. At the same time, there is a substantial pathway from unsuccessful execution of generated code into the Edit-RunEdit loop, indicating that many students refine solutions through manual edits and check their correctness by running them as well. Overall, the diagram suggests that both prompting and code editing are actively interleaved in students’ problem-solving workflows, though their relative use varies.
Figure 2(b) provides a turn-based flow diagram of the first nine turns per session111In the figure, we aggregate data across all problem categories, since the diagrams show the same overall pattern across task difficulty levels., clarifying how students interleave prompting and editing actions over time. The diagram alternates between odd-numbered turns, which represent messaging or terminal events, and even-numbered turns, which group executions and edit actions into sets. Three insights stand out from the visualization. First, a substantial number of sessions reach a solution within the first few turns, reflecting that initial prompts can often be effective. Second, manual editing typically appears as a reaction to unsuccessful executions of GenAI-generated code, as indicated by the thicker flow through the combined state {RunAI, RunEdit}, compared to the narrower {RunEdit}-only path. This confirms students mostly introduce manual edits after testing GenAI code. Finally, the ratio between the {RunAI}-only and {RunAI, RunEdit} flows remains relatively stable across turns, indicating that students introduce edits at a wide range of points during their sessions.
To complement these behavioral patterns, we analyzed responses to a Likert-style item asking what combination of prompting and editing they found most effective. They chose among five options (student share in brackets): 1. “Only natural language prompting” (5.36%); 2. “Mostly natural language prompting with a little manual code editing” (24.67%); 3. “An equal mix of natural language prompting and manual code editing” (44.77%); 4. “Mostly manual code editing with a little natural language prompting” (22.52%); 5. “Only manual code editing” (2.68%). Responses followed a normal distribution centered on an equal mix of prompting and editing, suggesting that students view the two approaches as complementary.
Furthermore, students responded to an open-ended question describing how they chose between prompting and editing. We found three key themes: prompting for speed, editing for precision, and learning with AI. Many students emphasized the efficiency of using prompting to quickly generate initial solutions, turning to manual edits for small fixes or refinements, as one student noted: “I used AI prompts first to get the basic code quickly. When I needed small fixes, I edited the code myself.” Others described a hybrid strategy in which AI-generated code served as a starting point that they incrementally corrected or refined through manual editing to ensure accuracy and handle edge cases. Finally, several students framed their choices in terms of learning, using prompting to explore unfamiliar concepts and editing to consolidate understanding and retain control over the solution. As one student reflected, “If it is a question I know how to do, then I use manual code, otherwise I use a prompt and learn from the AI generated code.”
4.2. RQ2: Effects of Task Complexity
To study how editing behavior varies with task difficulty, we analyze interaction logs across problem categories. For our analysis, we use the final code version from each Edit-RunEdit loop and look at four indicators of editing behavior in Figure 3.
Figures 3(a) and 3(b) show that editing was more common when students worked on more complex problems. In Figure 3(a), only about of students edited at least one Basic Functions problem, compared to roughly in Advanced Functions and in Classes & Functions. Figure 3(b) shows the same trend at the session level, with only about of Basic Functions sessions containing any edit, compared to around in Advanced Functions and in Classes & Functions (Basic vs. Advanced: ; Basic vs. Classes & Functions: ).222For pairwise comparisons between the three problem categories, we applied Bonferroni correction (Dunn, 1959) with a factor of . Because Advanced Functions and Classes & Functions were optional, we repeated the analysis on the students who attempted at least one problem in each category (visualizations not shown due to space constraints). The overall trends remained consistent.
Figure 3(c) examines the nature of edits, which remains stable across categories. More concretely, most edit cycles involved a single-line change, and the number of edited lines did not significantly differ by task difficulty. In addition, edits were typically followed by only a few additional interaction steps before the session finished (Figure 3(d)). This suggests that regardless of problem complexity, students primarily used editing for small, corrective adjustments rather than extensive rewrites.
4.3. RQ3: Effects of Programming Competence
To understand how programming competence relates to editing behavior, we analyze Basic Functions, the only mandatory category. This choice limits confounding effects from task difficulty and optional participation. We measure competence using each student’s mean score across the three invigilated code-writing tests used for course assessment (see Section 3). We split students into two groups, the HighScore group (mean ) and the LowScore group (mean ), with Figure 4 comparing their code-editing behavior.
Figures 4(a) and 4(b) show that students in the HighScore group were less likely to use manual editing. In Figure 4(a), across all problems (i.e., Overall in the figure), only about of sessions from the HighScore group contained any Edit action, compared to roughly of sessions from the LowScore group (). Figure 4(b) shows the same pattern for RunEdit action, with about of sessions from the HighScore group containing any RunEdit action across all problems (i.e., Overall in the figure), compared to around for the LowScore group (). We then examined whether this difference is specific to editing or reflects lower overall interaction. Figure 4(c) shows that students in the HighScore group solved the problems quicker, requiring significantly fewer messages per session before success ().
Together, these results suggest that higher-performing students reached correct solutions with fewer iterations across the full workflow, and therefore entered the edit and re-run loop less often. One possible explanation is that these students were able to think through the problem more effectively and formulate more accurate prompts early on, reducing the need for later manual edits after unsuccessful runs. In fact, prior work indicates that effective GenAI use relies on robust metacognitive skills, such as planning and monitoring one’s problem-solving process (Tankelevitch et al., 2024); overall, these results align with evidence that students with higher scores and confidence may use GenAI more effectively and experience fewer difficulties (Prather et al., 2024b).
4.4. Limitations
We acknowledge several limitations of our study that could be tackled in future work. First, only the Basic Functions problems were mandatory, while more complex tasks were optional, introducing potential self-selection bias. Future studies could mitigate this by assigning tasks of varying difficulty to all students. Second, we did not conduct an A/B study to compare prompting-only and prompting-plus-editing conditions in a controlled way. A randomized study where editing capabilities are selectively enabled could more directly measure the benefits of code editing. Third, we did not have student demographic data or their prior experience on using GenAI available to us, which limited the type of analyses we could do. Fourth, as we used a custom-built tool, it is possible the findings might not generalize to standard GenAI programming tools. Finally, our analysis spans only edited code that was either executed or sent as an attachment, excluding how much time or effort students spent editing. Analyzing intermediate drafts or more granular timestamps could provide deeper insights into students’ editing behavior.
5. Conclusion
Our study investigated a key question in AI-assisted programming: when do students revise their prompt, and when do they dive into the code? We found that students combined prompting and manual editing in their workflow, typically introducing edits after a generated solution failed execution. These manual edits were usually concise, serving as targeted, last-mile corrections. Student reflections highlighted prompting for efficiency, editing for precision, and both as valuable tools for learning. Edits became more common as tasks increased in complexity, though higher-performing students relied more consistently on prompting alone, reflecting their ability to craft effective prompts. These findings emphasize the need to support both skills in instruction, teaching students how to decide when to re-prompt versus apply an edit after testing. Tools and teaching materials can also offer scaffolding around the edit–run loop when students struggle. Developing this flexibility requires a solid understanding of the problem and the underlying programming concepts.
We believe our work opens up exciting ideas for future research. Future work should explore how to help students craft precise prompts while building fluency in concise, high-impact edits. First, studies could introduce controlled bugs into otherwise correct completions to cultivate students’ ability to detect and correct subtle issues. Second, research might investigate how students make strategic decisions between prompting and editing by presenting tasks that require them to justify which approach would be more effective in a given context. Third, longer-term investigations could track how editing behaviors evolve in more open-ended domains such as data analysis, visualization, or machine learning, where last-mile fixes may grow into broader refactors, offering deeper insight into how students balance prompting with manual control of code.
Acknowledgements.
Funded/Cofunded by the European Union (ERC, TOPS, 101039090). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.References
- (1)
- Acun and Acun (2023) Cagla Acun and Ramazan Acun. 2023. GAI-Enhanced Assignment Framework: A Case Study on Generative AI Powered History Education. In NeurIPS’23 Workshop on Generative AI for Education.
- Amoozadeh et al. (2024) Matin Amoozadeh, Daye Nam, Daniel Prol, Ali Alfageeh, James Prather, Michael Hilton, Sruti Srinivasa Ragavan, and Amin Alipour. 2024. Student-AI Interaction: A Case Study of CS1 Students. In Proceedings of the Koli Calling International Conference on Computing Education Research (Koli Calling).
- Becker et al. (2019) Brett A. Becker, Paul Denny, Raymond Pettit, Durell Bouchard, Dennis J. Bouvier, Brian Harrington, Amir Kamil, Amey Karkare, Chris McDonald, Peter-Michael Osera, Janice L. Pearce, and James Prather. 2019. Compiler Error Messages Considered Unhelpful: The Landscape of Text-Based Programming Error Message Research. In Proceedings of the Working Group Reports of the Conference on Innovation and Technology in Computer Science Education (ITiCSE).
- Cambaz and Zhang (2024) Doga Cambaz and Xiaoling Zhang. 2024. Use of AI-driven Code Generation Models in Teaching and Learning Programming: a Systematic Literature Review. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Denny et al. (2024a) Paul Denny, Sumit Gulwani, Neil T. Heffernan, Tanja Käser, Steven Moore, Anna N. Rafferty, and Adish Singla. 2024a. Generative AI for Education (GAIED): Advances, Opportunities, and Challenges. CoRR abs/2402.01580 (2024).
- Denny et al. (2024b) Paul Denny, Juho Leinonen, James Prather, Andrew Luxton-Reilly, Thezyrie Amarouche, Brett A. Becker, and Brent N. Reeves. 2024b. Prompt Problems: A New Programming Exercise for the Generative AI Era. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Denny et al. (2011) Paul Denny, Andrew Luxton-Reilly, Ewan D. Tempero, and Jacob Hendrickx. 2011. Understanding the syntax barrier for novices. In Proceedings of the Innovation and Technology in Computer Science Education Conference (ITiCSE).
- Denny et al. (2024c) Paul Denny, Stephen MacNeil, Jaromir Savelka, Leo Porter, and Andrew Luxton-Reilly. 2024c. Desirable Characteristics for AI Teaching Assistants in Programming Education. In Proceedings Companion of the Conference on Innovation and Technology in Computer Science Education (ITiCSE).
- Doughty et al. (2024) Jacob Doughty, Zipiao Wan, Anishka Bompelli, Jubahed Qayum, Taozhi Wang, Juran Zhang, Yujia Zheng, Aidan Doyle, Pragnya Sridhar, Arav Agarwal, Christopher Bogart, Eric Keylor, Can Kültür, Jaromír Savelka, and Majd Sakr. 2024. A Comparative Study of AI-Generated (GPT-4) and Human-crafted MCQs in Programming Education. In Proceedings of the Australasian Computing Education Conference (ACE).
- Dunn (1959) Olive Jean Dunn. 1959. Estimation of the Medians for Dependent Variables. The Annals of Mathematical Statistics (1959).
- Edwards et al. (2020) John Edwards, Joseph Ditton, Dragan Trninic, Hillary Swanson, Shelsey Sullivan, and Chad D. Mano. 2020. Syntax Exercises in CS1. In Proceedings of the International Computing Education Research Conference (ICER).
- Hüttel (2025) Hans Hüttel. 2025. On Program Synthesis and Large Language Models. Commun. ACM (2025).
- Kazemitabaar et al. (2023) Majeed Kazemitabaar, Xinying Hou, Austin Z. Henley, Barbara Jane Ericson, David Weintrop, and Tovi Grossman. 2023. How Novices Use LLM-based Code Generators to Solve CS1 Coding Tasks in a Self-Paced Learning Environment. In Proceedings of the Koli Calling International Conference on Computing Education Research (Koli Calling).
- Kazemitabaar et al. (2024) Majeed Kazemitabaar, Runlong Ye, Xiaoning Wang, Austin Zachary Henley, Paul Denny, Michelle Craig, and Tovi Grossman. 2024. CodeAid: Evaluating a Classroom Deployment of an LLM-based Programming Assistant that Balances Student and Educator Needs. In Proceedings of the Conference on Human Factors in Computing Systems (CHI).
- Kerslake et al. (2024) Chris Kerslake, Paul Denny, David H. Smith, James Prather, Juho Leinonen, Andrew Luxton-Reilly, and Stephen MacNeil. 2024. Integrating Natural Language Prompting Tasks in Introductory Programming Courses. In Proceedings of the Virtual Global Computing Education Conference (SIGCSE Virtual).
- Leinonen et al. (2023) Juho Leinonen, Paul Denny, Stephen MacNeil, Sami Sarsa, Seth Bernstein, Joanne Kim, Andrew Tran, and Arto Hellas. 2023. Comparing Code Explanations Created by Students and Large Language Models. In Proceedings of the Conference on Innovation and Technology in Computer Science Education (ITiCSE).
- Ma et al. (2024) Qianou Ma, Hua Shen, Kenneth Koedinger, and Sherry Tongshuang Wu. 2024. How to Teach Programming in the AI Era? Using LLMs as a Teachable Agent for Debugging. In Proceedings of the International Conference on Artificial Intelligence in Education (AIED).
- MacNeil et al. (2023) Stephen MacNeil, Andrew Tran, Arto Hellas, Joanne Kim, Sami Sarsa, Paul Denny, Seth Bernstein, and Juho Leinonen. 2023. Experiences from Using Code Explanations Generated by Large Language Models in a Web Software Development E-Book. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Mann and Whitney (1947) Henry B Mann and Donald R Whitney. 1947. On a Test of Whether One of Two Random Variables is Stochastically Larger than the Other. The Annals of Mathematical Statistics (1947).
- Markel et al. (2023) Julia M. Markel, Steven G. Opferman, James A. Landay, and Chris Piech. 2023. GPTeach: Interactive TA Training with GPT-based Students. In Proceedings of the Conference on Learning @ Scale (L@S).
- Nguyen et al. (2024b) Manh Hung Nguyen, Sebastian Tschiatschek, and Adish Singla. 2024b. Large Language Models for In-Context Student Modeling: Synthesizing Student’s Behavior in Visual Programming from One-Shot Observation. In Proceedings of the International Conference on Educational Data Mining (EDM).
- Nguyen et al. (2025) Manh Hung Nguyen, Alkis Gotovos Victor-Alexandru Pădurean, Sebastian Tschiatschek, and Adish Singla. 2025. Synthesizing High-Quality Programming Tasks with LLM-based Expert and Student Agents. CoRR abs/2504.07655 (2025).
- Nguyen et al. (2024a) Sydney Nguyen, Hannah McLean Babe, Yangtian Zi, Arjun Guha, Carolyn Jane Anderson, and Molly Q. Feldman. 2024a. How Beginning Programmers and Code LLMs (Mis)read Each Other. In Proceedings of the Conference on Human Factors in Computing Systems (CHI).
- Oli et al. (2025) Priti Oli, Rabin Banjade, Andrew M. Olney, and Vasile Rus. 2025. Can LLMs Identify Gaps and Misconceptions in Students’ Code Explanations? CoRR abs/2501.10365 (2025).
- OpenAI (2024) OpenAI. 2024. GPT-4o mini: Advancing Cost-efficient Intelligence. https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/.
- Padurean et al. (2025) Victor-Alexandru Padurean, Paul Denny, Alkis Gotovos, and Adish Singla. 2025. Prompt Programming: A Platform for Dialogue-based Computational Problem Solving with Generative AI Models. In Proceedings of the Innovation and Technology in Computer Science Education Conference (ITiCSE).
- Pădurean et al. (2025) Victor-Alexandru Pădurean, Paul Denny, and Adish Singla. 2025. BugSpotter: Automated Generation of Code Debugging Exercises. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Pădurean et al. (2024) Victor-Alexandru Pădurean, Georgios Tzannetos, and Adish Singla. 2024. Neural Task Synthesis for Visual Programming. Transactions on Machine Learning Research (TMLR) (2024).
- Phung et al. (2023) Tung Phung, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares. 2023. Generating High-Precision Feedback for Programming Syntax Errors using Large Language Models. In Proceedings of the International Conference on Educational Data Mining (EDM).
- Prather et al. (2023) James Prather, Paul Denny, Juho Leinonen, Brett A. Becker, Ibrahim Albluwi, Michelle Craig, Hieke Keuning, Natalie Kiesler, Tobias Kohn, Andrew Luxton-Reilly, Stephen MacNeil, Andrew Petersen, Raymond Pettit, Brent N. Reeves, and Jaromír Savelka. 2023. The Robots Are Here: Navigating the Generative AI Revolution in Computing Education. In Proceedings of the Working Group Reports of the Conference on Innovation and Technology in Computer Science Education (ITiCSE).
- Prather et al. (2024a) James Prather, Brent N. Reeves, Paul Denny, Juho Leinonen, Stephen MacNeil, Andrew Luxton-Reilly, João Orvalho, Amin Alipour, Ali Alfageeh, Thezyrie Amarouche, Bailey Kimmel, Jared Wright, Musa Blake, and Gweneth Barbre. 2024a. Breaking the Programming Language Barrier: Multilingual Prompting to Empower Non-Native English Learners. CoRR abs/2412.12800 (2024).
- Prather et al. (2024b) James Prather, Brent N. Reeves, Juho Leinonen, Stephen MacNeil, Arisoa S. Randrianasolo, Brett A. Becker, Bailey Kimmel, Jared Wright, and Ben Briggs. 2024b. The Widening Gap: The Benefits and Harms of Generative AI for Novice Programmers. In Proceedings of the Conference on International Computing Education Research (ICER).
- Schmucker et al. (2024) Robin Schmucker, Meng Xia, Amos Azaria, and Tom M. Mitchell. 2024. Ruffle &Riley: Insights from Designing and Evaluating a Large Language Model-Based Conversational Tutoring System. In Proceedings of the International Conference on Artificial Intelligence in Education (AIED).
- Shah et al. (2025) Anshul Shah, Anya Chernova, Elena Tomson, Leo Porter, William G. Griswold, and Adalbert Gerald Soosai Raj. 2025. Students’ Use of GitHub Copilot for Working with Large Code Bases. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Tankelevitch et al. (2024) Lev Tankelevitch, Viktor Kewenig, Auste Simkute, Ava Elizabeth Scott, Advait Sarkar, Abigail Sellen, and Sean Rintel. 2024. The Metacognitive Demands and Opportunities of Generative AI. In Proceedings of the Conference on Human Factors in Computing Systems (CHI).
- Vadaparty et al. (2024) Annapurna Vadaparty, Daniel Zingaro, David H. Smith IV, Mounika Padala, Christine Alvarado, Jamie Gorson Benario, and Leo Porter. 2024. CS1-LLM: Integrating LLMs into CS1 Instruction. In Proceedings of the Innovation and Technology in Computer Science Education Conference (ITiCSE).
- Wang et al. (2024) Sierra Wang, John C. Mitchell, and Chris Piech. 2024. A Large Scale RCT on Effective Error Messages in CS1. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Yellin (2023) Daniel M. Yellin. 2023. The Premature Obituary of Programming. Commun. ACM (2023).
- Zamfirescu-Pereira et al. (2023) J.D. Zamfirescu-Pereira, Laryn Qi, Bjorn Hartmann, John DeNero, and Narges Norouzi. 2023. Conversational Programming with LLM-Powered Interactive Support in an Introductory Computer Science Course. In NeurIPS’23 Workshop on Generative AI for Education.
- Zhang et al. (2024) Jialu Zhang, José Pablo Cambronero, Sumit Gulwani, Vu Le, Ruzica Piskac, Gustavo Soares, and Gust Verbruggen. 2024. PyDex: Repairing Bugs in Introductory Python Assignments using LLMs. Proceedings of the ACM on Programming Languages (2024).